Prepare any necessary journal entries for prior period adjustments based on the below information. Number each entry #1-9.

Polaris Corporation purchases ATVs from manufacturers and sells them to customers. Polaris had the following income-related information for the fiscal year 20Y4:

1. Polaris had sales of $1,500,000 during the year.

2. At the beginning of 20Y4, Polaris changed its inventory costing assumption from the average cost method to the FIFO method. If Polaris had used the FIFO method in prior years, 20Y3 cost of goods sold would have been lower by $30,000.

3. Polaris uses the periodic inventory system, and the newly adopted FIFO cost-flow method will be used to calculate cost of goods sold for 20Y4. After considering the change to FIFO discussed in item #3, Polaris had 15 ATVs in inventory at the start of 20Y4 with a cost of $7,000 each. During the year, Polaris purchased 95 ATVs at a cost of $8,000 each. A physical count indicated that there were 10 ATVs in inventory at the end of 20Y4.

4. Polaris had selling, general, and administrative expenses of $340,000 during 20Y4.

5. At the beginning of 20Y2, Polaris purchased equipment for $62,000 that had an estimated useful life of 7 years and an estimated salvage value of $6,000. However, the bookkeeper made an error and used a 10-year useful life in computing depreciation for years 20Y2 and 20Y3. Depreciation for 20Y4 was calculated correctly and is included in SG&A expenses in #4.

6. One of Polaris’s warehouses was destroyed in a fire during 20Y4. At the time of the fire, the warehouse had a book value of $200,000 and accumulated depreciation of $90,000. Polaris did not receive insurance proceeds related to the warehouse fire.

7. The company disposed of its Southwest geographic division during 20Y4, and the disposal represented a strategic shift. The division generated a pre-tax loss of $100,000 from operations during the year, and Polaris recognized a $310,000 pre-tax gain on the disposal.

8. All income statement items are taxed at 30%
Answer #1

the journal entries for the polaris corporation as answered in attachment

1- first is sales entry

2- change in valuation of inventory mthod to fifo has led to increase in profit to $30000, so amount will be taxed @30% on $30000

3-gross profit calculated

4- selling expenses

5-last 2 year deprication wrongly calculated as it was charged @5600 but it should be charged @8000 so the last 2 year deprication is calculated and adjusted in current year (8000-5600)= 2400*2=4800